# Forecasting mortality rates with functional signatures

### v1

@@ Abstract
This study introduces an innovative methodology for mortality forecasting, which integrates signature-based methods within the functional data framework of the Hyndman-Ullah (HU) model. This new approach, termed the Hyndman-Ullah with truncated signatures (HUts) model, aims to enhance the accuracy and robustness of mortality predictions. By utilizing signature regression, the HUts model aims to capture complex, nonlinear dependencies in mortality data which enhances forecasting accuracy across various demographic conditions. The model is applied to mortality data from 12 countries, comparing its forecasting performance against classical models like the Lee-Carter model and variants of the HU models across multiple forecast horizons. Our findings indicate that overall the HUts model not only provides more precise point forecasts but also shows robustness against data irregularities, such as those observed in countries with historical outliers. The integration of signature-based methods enables the HUts model to capture complex patterns in mortality data, making it a powerful tool for actuaries and demographers. Prediction intervals are also constructed using bootstrapping methods.
@@ 1 Introduction
Mortality modeling and forecasting are quintessential to the disciplines of actuarial science, demography, and public health, serving as critical tools for socio-economic planning and risk management. The primary objective of mortality forecasts is to provide reliable estimates of future mortality rates and life expectancies, which shape various policy and financial strategies. Understanding and anticipating mortality trends is fundamental for government agencies to adequately plan and allocate resources in healthcare, social security, and public welfare sectors. As populations age and demographic shifts occur, precise mortality forecasts are indispensable for sustainable public policy and healthcare system planning to guide governments in making informed decisions about retirement ages, healthcare funding, and pensions, ensuring that social security systems remain robust in the face of demographic changes.
In the insurance and pension industries, mortality modeling takes part in the pricing of life insurance and annuities, as well as the management of pension funds. Actuaries rely on these models to calculate premiums, reserves, and pension contributions that align with predicted mortality rates, thus safeguarding the financial health of these institutions against longevity risks (Deprez et al., 2017). Moreover, accurate mortality projections are critical in the context of global challenges such as the increasing life expectancy, coupled with declining fertility rates in many parts of the world, poses additional challenges with the shifting age structure of populations. These demographic trends amplify the importance of mortality forecasting, as they have profound implications for the economic sustainability of pension systems and public health policies (Bjerre, 2022). Thus, the practice of mortality modeling and forecasting not only supports financial and actuarial operations but also facilitates broader strategic planning and policy-making that accommodate the evolving demographic landscapes, ultimately ensuring the economic and social stability of societies.
@@ 1.1 Mortality modelling and forecasting
The seminal work of Lee and Carter (1992) was undoubtedly one of the most influential extrapolative models to be introduced to forecasts mortality rates. The Lee-Carter (LC) model
$\ln{(m_{x,t})}=a_{x}+b_{x}k_{t}+e_{x,t},$
@@
is an age-period mortality model with a log-bilinear structure that models the logarithm of central death rates, $m_{x,t}$ with a location parameter $a_{x}$ that accounts for the average logarithm of rates across the specified time period, a singular value decomposition derived bilinear term that captures the age pattern and mortality trends with a term that reflects the rate of mortality change of a particular age, $b_{x}$, and a mortality index that varies over time, $k_{t}$, and the error term $e_{x,t}$. Forecasts of the LC model are then produced by extrapolating the mortality index with time series methods where in most cases a random walk with drift would suffice. While the LC model was initially applauded for its simple structure and straightforward implementation, it did have its fair share of limitations which were subsequently addressed in various extensions of the LC model (Lee and Miller, 2001; Booth et al., 2002). Extensions were also not limited to adjustments to the original LC model as seen in the deployment of generalised linear models (Renshaw and Haberman, 2003) and Bayesian statistics (Czado et al., 2005). Readers are directed to Basellini et al. (2023) for a more recent review on LC methods since the introduction of the LC model.
@@
The LC extension of interest in this paper involves the functional data paradigm (Ramsay and Silverman, 2005). The Hyndman-Ullah (HU) model (Hyndman and Ullah, 2007) is a generalisation of the LC model that uses concepts from functional data analysis to better model the mortality rates by exploring the variations within the data from a functional perspective. The use of nonparametric smoothing techniques on the age-specific mortality rates allowed for the removal of the LC model’s homoskedasticity assumption and realise an underlying smooth curve that can be decomposed using functional principal component analysis with their orthonormal basis. Component wise it is similar to the LC model except it allows for multiple principal components to be included, accounting more variation into the model which Hyndman and Ullah (2007) noted is the reason for its superior performance for longer-term forecast. Comparisons also show that the HU model outperformed the LC, Lee-Miller, and Booth-Maindonald-Smith models. Recent extensions of the HU model have been for the modelling of subpopulations (Jiménez-Varón et al., 2024; Shang et al., 2022; Shang and Hyndman, 2017). In light of the substantial contributions made by the HU model, there is still a growing interest to explore new modifications to improve its efficacy and adaptability. This paper explores the integration of signatures into the HU model, to better capture the nonlinear patterns often present in mortality data of a single population. This approach seeks to merge established statistical techniques with innovative mathematical insights, potentially enriching the analytical landscape of actuarial science.
@@ 1.2 Signatures from rough path theory
The signature is a sophisticated mathematical tool that is capable of summarising complex, high-dimensional data paths through their iterated integrals in an informative and compact package as a power series that contains the tensor coefficients. Originally introduced by Chen (1957) and further developed in the context of rough path theory by Lyons et al. (2007), they have found extensive applications across multiple fields, particularly where the analysis of dynamic systems is crucial. Their effectiveness has been demonstrated in various applications such as sepsis detection (Morrill et al., 2020), identifying Chinese handwriting (Xie et al., 2018), and in diagnosing psychological disorders (Wang et al., 2020). The increasing interest in signatures stems primarily from their ability to effectively condense information from complex sequences without losing critical details, making them exceptionally useful for detection work. This efficiency in capturing the essence of data paths while minimizing complexity is what drives their growing popularity, especially in areas dealing with large volumes of sequential data.
@@
Signature regression is further highlighted through various applications across different fields. For instance, Levin et al. (2016) utilized the universal properties of signatures to characterize functional relationships within datasets, leading to the development of the expected signature model. This model leverages linear regression on signature features to summarize the conditional distribution of responses effectively. It has demonstrated superior forecasting abilities compared to autoregressive and Gaussian Process models, especially in scenarios involving large datasets, thereby highlighting its efficiency in data stream summarization and significant dimension reduction. Building on these foundations, Fermanian (2022) introduced the signature linear model within a functional regression framework. This model operates by mapping functions to their signatures and employing ridge regression for prediction, showing enhanced capabilities in handling multidimensional data. By surpassing established methods like functional linear models and functional principal component regression, the signature linear model eliminates the need for certain assumptions typically required within functional settings, streamlining the modeling process.
@@
In the field of actuarial science, the use of signatures has started to make inroads, particularly in areas such as financial derivatives pricing (Lyons et al., 2020), risk management (Andrès et al., 2024), and optimal execution (Kalsi et al., 2020). However, the application of signatures in the life and mortality side of actuarial science remains less explored, indicating a gap for research and potential development. The introduction of signature regression into mortality forecasting presents a novel methodology that could improve the accuracy of predictions by leveraging the universality of signatures to effectively model complex, nonlinear relationships inherent in mortality data. It is therefore the aim of this article to introduce a functional mortality model by which is decomposed using truncated signatures in place of FPCA that also utilises the extrapolation method to produce forecasts. The model will from here on be referred to as the HU with truncated signatures (HUts) model and will be used to obtain point and interval forecasts of log mortality rates of several countries that are readily available from the Human Mortality Database (2024).
@@ 2.1 The Hyndman-Ullah (HU) model
The HU model is a generalised extension of the Lee-Carter model, employing the functional data paradigm (Ramsay and Silverman, 2005) and nonparametric smoothing to improve the forecasting of mortality rates. This model is distinguished by its use of multiple principal components, capturing a broader variation within the data, and applying more advanced time series forecasting methods than the random walk with drift model typically used in the LC model (Hyndman and Ullah, 2007).
@@
where $\epsilon_{t}$ is the standard normal error term and $\sigma_{t}(x)$ denotes the observational standard deviation which is allowed to vary with age (Hyndman and Ullah, 2007). This lets the HU model handle data heterogeneity and ensures demographic consistency by applying monocity constraints when using penalized regression splines to smooth the mortality curves (Hyndman and Ullah, 2007).
@@
Here, $\mu(x)$ represents the mean function, $\phi_{k}(x)$ are orthonormal basis functions, $\beta_{t,k}$ are time dependent coefficients, and $e_{t}(x)$ is the error function. Forecasts of mortality rates are done through the time dependent coefficients, $\beta_{t,k}$ which are fitted into univariate robust ARIMA models and extrapolated $h$-steps ahead.
A critical consideration in applying FPCA is its sensitivity to outliers. Extreme values can disproportionately influence the principal components derived from the data. This sensitivity may lead to biased or misleading results, especially when modeling phenomena such as mortality rates where spikes might occur due to extraordinary events like epidemics or wars. To address these issues, various adaptations of the HU model have been put forth. These include the robust HU (HUrob) model with robust FPCA (Hyndman and Ullah, 2007), which seeks to minimize the impact of outliers by using the $L_{1}$-median as the location measure, paired with the reflection-based algorithm for principal components analysis (Hubert et al., 2002) for a robust set of principal components; the weighted HU (wHU) model with weighted FPCA (Hyndman and Shang, 2009), which adjusts the influence of data points by placing more weight on recent data with the measure of location being a weighted average using geometrically decreasing weights. These modifications aim to enhance the stability and reliability of the model under conditions where data anomalies are present. Having the same goal in mind to make the model less sensitive to outliers, we extend the literature by exploring the use of signature regression on the signature of age specific mortality rates as an alternative to enhancing the robustness of mortality forecasts.
@@ 2.2.1 The concept of signatures
Definition 2.1 (Signature of a path (Fermanian, 2022))
@@
which contains $\frac{d^{m+1}-1}{d-1}$ number of signature coefficients. Note that the number of signature coefficients grows exponentially with the truncation order.
@@ 2.3 Embedding with signatures
In the proceeding sectionss, we will use these embedding transformation to prepare our data paths before computing its signature so as to enhance the signatures’ ability to capture and reflect the complex dynamics of the sequences.
@@ 2.4 Signature regression
Definition 2.2 (Universal nonlinearity (Bonnier et al., 2019))
@@ 2.5 The Hyndman-Ullah model with truncated signatures
We are now prepared to integrate signatures into the Hyndman-Ullah (HU) model, which we will refer to as the HU with truncated signatures (HUts) model. The integration of signatures with functional data models has been documented by Fermanian (2022) and Frévent (2023). However, these studies primarily applied signature regression in a functional regression framework with scalar responses and functional covariates, where signatures of the functional covariates were used directly. In contrast, the HU model operates more like a functional regression model with functional responses, making the direct application of signature regression more complex. We plan to consider using the signature of cross-sectional data from these functional covariates to apply signature regression at each point throughout the functional domain and then reconstruct a functional response. The HUts model is presented as follows:
@@
Consider the time series of the smoothed mortality rates for age $x_{i}$, $\{\hat{f}_{t}(x_{i})\}_{t=1}^{n}$. Combining the basepoint augmentation, time augmentation, and the lead-lag transformation, the augmented path can be transformed as follows:
$(\displaystyle\tilde{X_{i}}=\Bigl{(}$$(t_{0},0,0),(t_{1},\hat{f}_{1}(x_{i}),\hat{f}_{1}(x_{i})),(t_{2},\hat{f}_{2}(x_{i}),\hat{f}_{1}(x_{i})),(t_{3},\hat{f}_{2}(x_{i}),\hat{f}_{2}(x_{i})),$
$(t_{4},\hat{f}_{3}(x_{i}),\hat{f}_{2}(x_{i})),\ldots,(t_{n},\hat{f}_{n}(x_{i}),\hat{f}_{n}(x_{i}))\Bigr{)},$
where $t$ is an ordered sequence within the interval $[0,1]$ where $t_{0}=0<t_{1}<t_{2}<\ldots<t_{n}=1$, marking the times of observations. This structured augmentation ensures that each data point $x_{i}$ is not only represented by its current value but also linked to its immediate past, enhancing the path’s dimensionality and capturing both the direction and history within the data, thereby enriching the data representation for signature computation. The transformed time series $\tilde{X}_{i}$ is then used to compute the truncated signature $S^{m}(\tilde{X}_{i})$.
@@
$S=\begin{bmatrix}S^{m}(\tilde{X}_{1})\\ \vdots\\ S^{m}(\tilde{X}_{q})\end{bmatrix}_{q\times s_{d}(m)},$
$Y_{t}=\begin{bmatrix}f^{*}_{t}(x_{1})\\ \vdots\\ f^{*}_{t}(x_{q})\end{bmatrix}_{q\times 1}.$
We then apply PCA to obtain the orthogonal principal component scores by singular value decomposition, $S=U\Sigma V^{T}$ where the first $K$ principal component scores are the first $K$ columns of $Z=U\Sigma$. Then, to obtain $\hat{\beta}_{t}$ is equivalent to solving
@@
where $Z_{K}$ is a $q\times K$ matrix where the $i^{th}$ row contains the normalised principal component of the truncated signature of order $m$ of age $x_{i}$,

### v2

@@ Abstract
This study introduces an innovative methodology for mortality forecasting, which integrates signature-based methods within the functional data framework of the Hyndman-Ullah (HU) model. This new approach, termed the Hyndman-Ullah with truncated signatures (HUts) model, aims to enhance the accuracy and robustness of mortality predictions. By utilizing signature regression, the HUts model is able to capture complex, nonlinear dependencies in mortality data which enhances forecasting accuracy across various demographic conditions. The model is applied to mortality data from 12 countries, comparing its forecasting performance against variants of the HU models across multiple forecast horizons. Our findings indicate that overall the HUts model not only provides more precise point forecasts but also shows robustness against data irregularities, such as those observed in countries with historical outliers. The integration of signature-based methods enables the HUts model to capture complex patterns in mortality data, making it a powerful tool for actuaries and demographers. Prediction intervals are also constructed with bootstrapping methods.
@@ 1 Introduction
Mortality modeling and forecasting are quintessential to the disciplines of actuarial science, demography, and public health, serving as critical tools for socio-economic planning and risk management. Their primary objective is to provide reliable estimates of future mortality rates and life expectancies, which are essential for shaping policy and financial strategies. For government agencies, anticipating mortality trends is fundamental for effective resource allocation in healthcare, social security, and public welfare sectors. Precise mortality forecasts enable informed decisions on retirement ages, healthcare funding, and pensions, ensuring that social security systems remain robust amid demographic changes.
In the insurance and pension industries, mortality modeling takes part in the pricing of life insurance, annuities and managing pension funds. Actuaries rely on these models to calculate premiums, reserves, and pension contributions that align with predicted mortality rates, thus safeguarding institutions against longevity risks (Deprez et al., 2017). Accurate mortality projections are also critical in addressing global challenges posed by increasing life expectancy and declining fertility rates which shift the age structure of populations. These demographic trends have profound implications for the economic sustainability of pension systems and public health policies (Bjerre, 2022). Consequently, the practice of mortality modeling and forecasting not only supports financial and actuarial operations but also facilitates broader strategic planning and policy-making to accommodate the evolving demographic landscapes, ultimately ensuring the economic and social stability of societies.
@@ 1.1 Mortality modelling and forecasting
The seminal work of Lee and Carter (1992) was undoubtedly one of the most influential extrapolative models to be introduced to forecast mortality rates. While the Lee-Carter (LC) model was initially applauded for its simple structure and straightforward implementation, it did have its fair share of limitations which were subsequently addressed in various extensions of the LC model (Lee and Miller, 2001; Booth et al., 2002). Extensions were also not limited to adjustments to the original LC model as seen in the deployment of generalised linear models (Renshaw and Haberman, 2003) and Bayesian statistics (Czado et al., 2005). Readers are directed to Basellini et al. (2023) for a more recent review of LC methods since the introduction of the LC model.
@@
The LC extension of interest in this paper involves the functional data paradigm (Ramsay and Silverman, 2005). The Hyndman-Ullah (HU) model (Hyndman and Ullah, 2007) is a generalisation of the LC model that uses concepts from functional data analysis to better model mortality rates by exploring variations within the data from a functional perspective. The use of nonparametric smoothing techniques on the age-specific mortality rates allowed for the removal of the LC model’s homoskedasticity assumption and realise an underlying smooth curve that can be decomposed using functional principal component analysis with their orthonormal basis. Component wise it is similar to the LC model except it allows for multiple principal components to be included, accounting more variation into the model which Hyndman and Ullah (2007) noted is the reason for its superior performance for longer-term forecast. Comparisons also show that the HU model outperformed the LC, Lee-Miller, and Booth-Maindonald-Smith models. Recent extensions of the HU model have been for the modelling of subpopulations (Jiménez-Varón et al., 2024; Shang et al., 2022; Shang and Hyndman, 2017). In light of the substantial contributions made by the HU model, there is still a growing interest to explore new modifications to improve its efficacy and adaptability.
With the rise of network approaches in modelling, there is growing interest in extending age-period models using deep learning techniques. Various aspects of the LC model have been coupled with deep learning applications, both in parameter estimation (Schnürch and Korn, 2022; Scognamiglio, 2022; Richman and Wüthrich, 2021; Hainaut, 2018; Miyata and Matsuyama, 2022) and in forecasting procedures (Perla et al., 2021; Bjerre, 2022; Nigri et al., 2019; Marino et al., 2023). Deep learning models fundamentally process streams of input data to approximate relationships. While they perform well, they can lack replicability and often require increased computational resources. Signatures offer a potential solution to these limitations as they too possess universality to approximate relationships arbitrarily well, reproducible, and efficient. Therefore, this paper aims to integrate signatures that better capture the nonlinear patterns often present in single-population mortality data into the HU model’s functional data framework.
@@ 1.2 Signatures from rough path theory
The signature is a sophisticated mathematical tool that is capable of summarising complex, high-dimensional data paths through their iterated integrals in an informative and compact package as a power series that contains the tensor coefficients. Originally introduced by Chen (1957) and further developed in the context of rough path theory by Lyons et al. (2007), they have found extensive applications across multiple fields, particularly where the analysis of dynamic systems is crucial. Their effectiveness has been demonstrated in various applications such as sepsis detection (Morrill et al., 2020), identifying Chinese handwriting (Xie et al., 2018), and in diagnosing psychological disorders (Wang et al., 2020). The increasing interest in signatures stems primarily from their ability to effectively condense information from complex sequences without losing critical details, making them exceptionally useful for detection work. This efficiency in capturing the essence of data paths while minimizing complexity is what drives their growing popularity, especially in areas dealing with large volumes of sequential data such as in the field of quantitative finance (Lyons et al., 2020; Kalsi et al., 2020; Cuchiero and Möller, 2024; Jaber and Gérard, 2024).
@@
Signature regression is further highlighted through various applications across different fields. For instance, Levin et al. (2016) utilized the universal properties of signatures to characterize functional relationships within datasets, leading to the development of the expected signature model. This model leverages linear regression on signature features to summarize the conditional distribution of responses effectively. It has demonstrated superior forecasting abilities compared to autoregressive and Gaussian Process models, especially in scenarios involving large datasets, thereby highlighting its efficiency in data stream summarization and significant dimension reduction.
@@
Building on these insights, Fermanian (2022) was the first to apply the concept of signatures to functional data analysis by introducing the signature linear model within a functional regression framework. Fermanian (2022) noted that signatures are naturally adapted to vector-valued functions, unlike other functional methods, and require little assumption on the regularity of functions to encode nonlinear geometric information. This model operates by mapping functions to their signatures and employing ridge regression for prediction, showing enhanced capabilities in handling multidimensional data. By surpassing established methods like functional linear models and functional principal component regression, the signature linear model eliminates the need for certain assumptions typically required within functional settings, streamlining the modeling process. This lays the foundation that suggests that signature regression is a good candidate to extend the functional data framework of the HU model.
In the field of actuarial science, the use of signatures has started to make inroads, particularly in areas such as validating economic scenarios (Andrès et al., 2024), and option pricing (Cuchiero et al., 2023, 2024). However, the application of signatures in the life and mortality side of actuarial science remains less explored, indicating a gap for research and potential development. The introduction of signature regression into mortality forecasting presents a novel methodology that could improve the accuracy of predictions by leveraging the universality of signatures to effectively model complex, nonlinear relationships inherent in mortality data. It is therefore the aim of this article to introduce a functional mortality model by which is decomposed using truncated signatures in place of FPCA that also utilises the extrapolation method to produce forecasts. The model will from here on be referred to as the HU with truncated signatures (HUts) model and will be used to obtain point and interval forecasts of log mortality rates of several countries that are readily available from the Human Mortality Database (2024).
@@ 2.1 The Hyndman-Ullah (HU) model
The HU model is a generalised extension of the LC model, employing the functional data paradigm (Ramsay and Silverman, 2005) and nonparametric smoothing to improve the forecasting of mortality rates. This model is distinguished by its use of multiple principal components, capturing a broader variation within the data, and applying more advanced time series forecasting methods than the random walk with drift model typically used in the LC model (Hyndman and Ullah, 2007).
@@
where $\epsilon_{t}$ is the standard normal error term and $\sigma_{t}(x)$ denotes the observational standard deviation which is allowed to vary with age. This lets the HU model handle data heterogeneity and ensures demographic consistency by applying monocity constraints when using penalized regression splines to smooth the mortality curves (Hyndman and Ullah, 2007).
@@
Here, $\mu(x)$ represents the mean function, $\phi_{k}(x)$ are orthonormal basis functions, $\beta_{t,k}$ are time dependent coefficients, and $e_{t}(x)$ is the error function. Forecasts of mortality rates are done through the time dependent coefficients, $\beta_{t,k}$ which are fitted into univariate ARIMA models and extrapolated $h$-steps-ahead.
A critical consideration in applying FPCA is its sensitivity to outliers. Extreme values can disproportionately influence the principal components derived from the data. This sensitivity may lead to biased or misleading results, especially when modeling phenomena such as mortality rates where spikes might occur due to extraordinary events like epidemics or wars. To address these issues, various adaptations of the HU model have been put forth. These include the robust HU (HUrob) model with robust FPCA (Hyndman and Ullah, 2007), which seeks to minimize the impact of outliers by using the $L_{1}$-median as the location measure, paired with the reflection-based algorithm for principal components analysis (Hubert et al., 2002) for a robust set of principal components; the weighted HU (wHU) model with weighted FPCA (Hyndman and Shang, 2009), which adjusts the influence of data points by placing more weight on recent data with the measure of location being a weighted average using geometrically decreasing weights. These modifications aim to enhance the stability and reliability of the model under conditions where data anomalies are present. Having the same objective in mind to make the model less sensitive to outliers, we extend the literature by exploring the use of signature regression on the signature of age specific mortality rates as an alternative to enhancing the robustness of mortality forecasts.
@@ 2.2.1 The concept of signatures
Definition 2.1 (Signature of a path (Fermanian, 2022)).
@@
which contains $s_{d}(m)=\frac{d^{m+1}-1}{d-1}$ number of signature coefficients. Note that the number of signature coefficients grows exponentially with the truncation order.
@@ 2.3 Embedding with signatures
There are various embedding techniques that can be used to augment paths, each suited to specific applications based on the aspects of the data they capture. Different embedding choices highlight different features of the data, making them appropriate for particular tasks (Fermanian, 2021). In this paper, we focus mainly on three types of embeddings: basepoint augmentation, time augmentation, and the lead-lag transformation.
@@
In the proceeding sections, we will use these embedding transformation to prepare our data paths before computing its signature so as to enhance the signatures’ ability to capture and reflect the complex dynamics of the sequences.
@@ 2.4 Signature regression
Definition 2.2 (Universal nonlinearity (Bonnier et al., 2019)).
@@ 2.5 The Hyndman-Ullah model with truncated signatures
We are now prepared to integrate signatures into the Hyndman-Ullah (HU) model, which we will refer to as the HU with truncated signatures (HUts) model. The integration of signatures with functional data models have been documented by Fermanian (2022) and Frévent (2023). However, these studies primarily applied signature regression in a functional regression framework with scalar responses and functional covariates, where signatures of the functional covariates were used directly. In contrast, the HU model operates more like a functional regression model with functional responses, making the direct application of signature regression more complex. We plan to consider using the signature of cross-sectional data from these functional covariates to apply signature regression at each point throughout the functional domain and then reconstruct a functional response. The HUts model is presented as follows:
@@
Consider the time series of the centered smoothed mortality rates for age $x_{i}$, $\{f^{*}_{t}(x_{i})\}_{t=1}^{n}$. Combining the basepoint augmentation, time augmentation, and the lead-lag transformation, the augmented path can be transformed as follows:
$(\displaystyle\tilde{X_{i}}=\Bigl{(}$$(t_{0},0,0),(t_{1},f^{*}_{1}(x_{i}),f^{*}_{1}(x_{i})),(t_{2},f^{*}_{2}(x_{i}),f^{*}_{1}(x_{i})),(t_{3},f^{*}_{2}(x_{i}),f^{*}_{2}(x_{i})),$
$(t_{4},f^{*}_{3}(x_{i}),f^{*}_{2}(x_{i})),\ldots,(t_{n},f^{*}_{n}(x_{i}),f^{*}_{n}(x_{i}))\Bigr{)},$
where $t$ is an ordered sequence within the interval $[0,1]$ and $t_{0}=0<t_{1}<t_{2}<\ldots<t_{n}=1$, marking the times of observations. This structured augmentation ensures that each data point $x_{i}$ is not only represented by its current value but also linked to its immediate past, enhancing the path’s dimensionality and capturing both the direction and history within the data, thereby enriching the data representation for signature computation. The transformed time series $\tilde{X}_{i}$ is then used to compute the truncated signature $S^{m}(\tilde{X}_{i})$.
@@
$S=\begin{bmatrix}S^{m}(\tilde{X}_{1})\\ \vdots\\ S^{m}(\tilde{X}_{q})\end{bmatrix}_{q\times s_{d}(m)}\;,\hskip 28.45274ptY_{t}=\begin{bmatrix}f^{*}_{t}(x_{1}^{*})\\ \vdots\\ f^{*}_{t}(x_{q}^{*})\end{bmatrix}_{q\times 1}.$ (2.1)
We then normalise $S$ to obtain $S^{*}$ before applying PCA to obtain the orthogonal principal component scores by singular value decomposition, $S^{*}=U\Sigma V^{T}$ where the first $K$ principal component scores are the first $K$ columns of $Z=U\Sigma$. Then, to obtain $\hat{\beta}_{t}$ is equivalent to solving
@@
where $Z_{K}$ is a $q\times K$ matrix where the $i^{th}$ row contains the principal component of $S^{*}$ of order $m$ of age $x_{i}$,
@@
Remark 2.3.
As time augmentation leads to identical values for the signature coefficients $S^{(1)},S^{(1,1)},S^{(1,1,1)},\ldots$, along each row of $S$ in Equation (2.1), certain columns of $S$ become constant. During normalisation, these zero-variance columns are excluded from being centered and scaled, ensuring a well-defined normalized matrix $S^{*}$. Consequently, when performing PCA on $S^{*}$, one of the principal components is expected to be constant.